21) How many moles of NCl3 contain 2.55 × 1024 chlorine atoms?

A) 1.41 moles NCl3

B) 4.23 moles NCl3

C) 12.7 moles NCl3

D) 7.87 moles NCl3

E) 2.82 moles NCl3

Can someone please explain step by step how to get the correct answer witch is A.
Answer #1

1 mole of NCl3 contains 3 moles of chlorine atoms

3 moles of chlorine atoms in 1 mole of NCl3

3*6.023*10^23 chlorine atoms in 1 mole of NCl3

2.55*10^24 chlorine atoms   = 1*2.55*10^24/(3*6.023*10^23)   = 1.41moles of NCl3

A) 1.41 moles NCl3 >>>>answer